# Cross-Embodied Co-Design for Dexterous Hands

### v1

@@ Abstract
Dexterous manipulation is limited by both control and design, without consensus as to what makes manipulators best for performing dexterous tasks. This raises a fundamental challenge: how should we design and control robot manipulators that are optimized for dexterity? We present a co-design framework that learns task-specific hand morphology and complementary dexterous control policies. The framework supports 1) an expansive morphology search space including joint, finger, and palm generation, 2) scalable evaluation across the wide design space via morphology-conditioned cross-embodied control, and 3) real-world fabrication with accessible components. We evaluate the approach across multiple dexterous tasks, including in-hand rotation with simulation and real deployment. Our framework enables an end-to-end pipeline that can design, train, fabricate, and deploy a new robotic hand in under 24 hours. The full framework will be open-sourced and available on our website.
@@ 3 Method
Figure 2: Overview of our method. Stage 1: We first randomly sample embodiments, including morphology and degrees of freedom. We then pre-train a morphology-conditioned policy across embodiments. Stage 2: The design is generated using modular grammars, with assembly rules selected based on prior performance. The cross-embodied policy is then deployed to evaluate different designs in simulation. Stage 3: Best designs and fine tuned control are then manufactured and deployed on real hardware across unseen objects.
@@ 3.2 Architecture and Training
Sim-to-Real Transfer. The best-ranked design from search is fine-tuned with domain randomization over actuator characteristics, contact and friction parameters, and object poses to improve robustness (Andrychowicz et al., 2020; Tobin et al., 2017). For deployment, we train the controller as a blind policy, i.e., removing object-state inputs, to match onboard sensing and reduce sim–to-real mismatch. There is no camera or tactile feedback in the system. The cross embodied policy for the respective design family is fine tuned by loading the checkpoint and tuning weights for the stateless observations. Manufacturing the real robot hand follows the same programmatic grammar used in simulation: the best selected design graph is directly converted to modular hardware specifications with 3D-printed components and precisely matched actuator constraints. A mounting interface is added to the palm, each joint and links are directly 3D printed, and then assembled with Dynamixel actuators. After tuning PID on hardware, the blind policy is deployed on the physical system.
@@ 4 Finger 10 9
We show the time used to rotate a given object 360 degrees, and a subset of the results is shown in Table 3. Failure to rotate within one minute is denoted as a dash. The three fingered robot hand executed grasps which were able to generalize more effectively to unseen objects. The anthropomorphic and four finger robot hands were able to blindly rotate a small subset of objects - three out of seventeen tested objects, as predicted to be lower performing by our algorithm. Objects often became stuck between finger gaits. Across the same seventeen objects, the three fingered robot hand generalized substantially better generalization. Across all seventeen objects, only two failed with one two small between the robots bling gait to reach. The three fingered morphology generalized across irregular objects including pine cones, pentagons, and soft objects. The gait of the three fingered robot hand is three altering diagonals, allowing the object to be rotated with little displacement. Please see videos for details.

### v2

@@ Abstract
Dexterous manipulation is limited by both control and design, without consensus as to what makes manipulators best for performing dexterous tasks. This raises a fundamental challenge: how should we design and control robot manipulators that are optimized for dexterity? We present a co-design framework that learns task-specific hand morphology and complementary dexterous control policies. The framework supports 1) an expansive morphology search space including joint, finger, and palm generation, 2) scalable evaluation across the wide design space via morphology-conditioned cross-embodied control, and 3) real-world fabrication with accessible components. We evaluate the approach across multiple dexterous tasks, including in-hand rotation with simulation and real deployment. Our framework enables an end-to-end pipeline that can design, train, fabricate, and deploy a new robotic hand in under 24 hours. Results and build guide are available at: https://an-axolotl.github.io/co-design-for-dexterity.github.io/
@@ 3 Method
Figure 2: Overview of our method. Stage 1: We first randomly sample embodiments, including morphology and degrees of freedom. We then pre-train a morphology-conditioned policy across embodiments. Stage 2: The design is generated using modular grammars, with assembly rules selected based on prior performance. The cross-embodied policy is then deployed to evaluate different designs in simulation. Stage 3: Best designs and control pairs are then manufactured and deployed on real hardware across unseen objects.
@@ 3.2 Architecture and Training
Sim-to-Real Transfer. The best-ranked design from search is selected and tuned alongside domain randomization over actuator characteristics, contact and friction parameters, and object poses to improve robustness (Andrychowicz et al., 2020; Tobin et al., 2017). We deploy the design conditioned controller as a blind policy, i.e., removing object-state inputs, to match onboard sensing and reduce sim–to-real mismatch. There is no camera or tactile feedback in the system.
Manufacturing the real robot hand follows the same programmatic grammar used in simulation: the best selected design graph is directly converted to modular hardware specifications with 3D-printed components and precisely matched actuator constraints. A mounting interface is added to the palm, each joint and links are directly 3D printed, and then assembled with Dynamixel actuators. After tuning PID on hardware, the blind policy is deployed on the physical system.
@@ 4 Finger 10 9
We show the time used to rotate a given object 360 degrees, and a subset of the results is shown in Table 3. Failure to rotate within one minute is denoted as a dash. The three fingered robot hand executed grasps which were able to generalize more effectively to unseen objects. The anthropomorphic and four finger robot hands were able to blindly rotate a small subset of objects - three out of seventeen tested objects, as predicted to be lower performing by our algorithm. Objects often became stuck between finger gaits. Across the same seventeen objects, the three fingered robot hand generalized substantially better generalization. Across all seventeen objects, only two failed with one two small between the robots bling gait to reach. The three fingered morphology generalized across irregular objects including pine cones, pentagons, and soft objects. The gait of the three fingered robot hand is three altering diagonals, allowing the object to be rotated with little displacement. Cross embodied policies also lead to improved generalization and dexterity over PPO trained for a single hand (Appendix 9.1.2). Please see video and website for details.